# MassSpecGym: A benchmark for the discovery and identification of molecules

### v1

@@ 2 Related work
Figure 4: Our MCES-based data split resolves the data-leakage issue abundant in prior work. The standard approach separates molecules with identical planar structures (2D InChIKeys) into different folds, disregarding minor molecular modifications. This leads to near-duplicate test molecules (with Tanimoto similarity > 0.85) being leaked from the training data (red), as shown in the example below. In contrast, our approach minimizes molecular edit distance (MCES) between training and test sets, ensuring distinct data folds (green).
@@ 3.2 Definition of the challenges
Cosine similarity between two spectra ranges from 0 to 1, where 1 corresponds to a perfect prediction. A key application of accurately predicting spectra from molecules is in molecular retrieval [55, 21, 23]. Accurate and scalable models enable the automatic annotation of molecular databases, bolstering the coverage of existing spectral libraries. Therefore, we can additionally use an analagous setup as a downstream task to evaluate spectrum predictions. Similarly to the molecule retrieval task defined previously, for a molecule-spectrum pair $(G,X)$, the set of candidates $\mathcal{C}$ comprises of $G$ and the set of molecules from a chemical database most similar to $G$. For each molecule $G_{i}\in\mathcal{C}$, we predict a spectrum $\hat{\mathbf{x}}_{i}$ and rank all candidates by decreasing cosine similarity between $\hat{\mathbf{x}}_{i}$ and $\mathbf{x}$. We evaluate the model by the rate at which $G$ is ranked in the top-$k$ hits in the sorted $\mathcal{C}$, using the same hit rate @ $k$ metrics as defined in the previous section.
@@ 4.1 Baseline models
The simplest baseline method for molecule retrieval, Random, sorts the candidate molecules $\mathcal{C}$ randomly. The second method, Fingerprint FFN, employs a feedforward neural network to predict the Morgan fingerprint of the target molecule. The candidates are then sorted based on their cosine similarity to the predicted fingerprint. Next, we evaluate MIST, a state-of-the-art deep learning approach, also based on fingerprint prediction. MIST assigns chemical subformulae to spectral peaks via energy-based modeling [22], then predicts a molecular fingerprint via a chemical formula-based transformer, and finally ranks the candidates by cosine similarity between the fingerprints [19]. Finally, we evaluate DeepSets [71]. The model processes spectra as sets of raw 2D peak representations, providing a complementary approach to FingerprintFFN and the state-of-the-art MIST which are based on fingerprint prediction. DeepSets + Fourier features enhances DeepSets by using Fourier features enabling more accurate modeling of m/z values [20].
@@
[] Cosine Similarity $\uparrow$ Jensen-Shannon Similarity $\uparrow$ Hit Rate @ 1 $\uparrow$ Hit Rate @ 5 $\uparrow$ Hit Rate @ 20 $\uparrow$ Precursor m/z 0.15 (0.14-0.17) 0.41 (0.40-0.42) 0.38 (0.21-0.62) 1.72 (1.32-2.18) 7.17 (6.32-8.04) FFN Fingerprint 0.25 (0.24-0.26) 0.48 (0.47-0.48) 8.44 (7.56-9.34) 21.43 (20.10-22.79) 38.57 (36.99-40.23) GNN 0.19 (0.18-0.20) 0.44 (0.44-0.45) 3.95 (3.37-4.62) 11.92 (10.87-13.00) 26.27 (24.83-27.82) FraGNNet 0.52 (0.51-0.53) 0.63 (0.63-0.64) 46.64 (44.98-48.26) 72.56 (71.18-74.00) 83.58 (82.34-84.75) Bonus chemical formulae challenge Precursor m/z - - 2.09 (1.66-2.59) 8.52 (7.65-9.53) 22.65 (21.26-24.01) FFN Fingerprint - - 7.62 (6.77-8.54) 22.70 (21.32-24.12) 44.12 (42.51-45.75) GNN - - 3.63 (3.05-4.29) 13.55 (12.46-14.68) 33.77 (32.26-35.37) FraGNNet - - 31.93 (30.40-33.50) 63.20 (61.64-64.76) 82.70 (81.45-83.93)
@@ 5 Conclusions
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90140). This work was also co-funded by the European Union (ERC, FRONTIER, 101097822). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. SH and AK are funded through Award R35GM148219 by the NIGMS of the National Institutes of Health. TP was supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon 2020 research and innovation programme under Marie Skłodowska-Curie grant agreement No. 891397. AY was supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) scholarship and a Vector Institute research grant. LZ was supported by NSERC. FW was supported by Alberta Machine Intelligence Institute (AMII), and NSERC grant. RG was was supported by NSERC, AMII. DSW was supported by NSERC, and Genome Canada, Genome British Columbia and Genome Alberta. HR was supported by NSERC and the Canada Research Chair (CRC) Program. FK, KD and SB are supported by Deutsche Forschungsgemeinschaft (BO 1910/23). FK and SB are supported by the Ministry for Economics, Sciences and Digital Society of Thuringia (Framework ProDigital, DigLeben-5575/10-9). NAH and SB are supported by the Thüringer Ministerium für Wirtschaft, Wissenschaft und Digitale Gesellschaft (TMWWDG) with funds from the European Union as part of the European Regional Development Fund (ERDF, 2023 VFE 0003). JH and WB acknowledge support by the University of Antwerp Research Fund. FH was supported by Deutsche Forschungsgemeinschaft (DFG, 528775510). LL was supported by NSF Award 2239869. CB was supported by the Czech Academy of Sciences PPLZ fellowship number L200552251.

### v2

@@ 2 Related work
Figure 4: Our MCES-based data split resolves the data-leakage issue abundant in prior work. The standard approach separates molecules with identical planar structures (2D InChIKeys) into different folds, disregarding minor molecular modifications. This leads to near-duplicate test molecules (with Tanimoto similarity > 0.85) being leaked from the training data (red), as shown in the example below. In contrast, our approach maximizes molecular edit distance (MCES) between training and test sets, ensuring distinct data folds (green).
@@ 3.2 Definition of the challenges
Cosine similarity between two spectra ranges from 0 to 1, where 1 corresponds to a perfect prediction. Jensen-Shannon similarity is reported as an additional metric (see Supplementary Information).
A key application of accurately predicting spectra from molecules is in molecular retrieval [55, 21, 23]. Accurate and scalable models enable the automatic annotation of molecular databases, bolstering the coverage of existing spectral libraries. Therefore, we can additionally use an analagous setup as a downstream task to evaluate spectrum predictions. Similarly to the molecule retrieval task defined previously, for a molecule-spectrum pair $(G,X)$, the set of candidates $\mathcal{C}$ comprises of $G$ and the set of molecules from a chemical database most similar to $G$. For each molecule $G_{i}\in\mathcal{C}$, we predict a spectrum $\hat{\mathbf{x}}_{i}$ and rank all candidates by decreasing cosine similarity between $\hat{\mathbf{x}}_{i}$ and $\mathbf{x}$. We evaluate the model by the rate at which $G$ is ranked in the top-$k$ hits in the sorted $\mathcal{C}$, using the same hit rate @ $k$ metrics as defined in the previous section.
@@ 4.1 Baseline models
The simplest baseline method for molecule retrieval, Random, sorts the candidate molecules $\mathcal{C}$ randomly. The second method, Fingerprint FFN, employs a feedforward neural network to predict the Morgan fingerprint of the target molecule. The candidates are then sorted based on their cosine similarity to the predicted fingerprint. Next, we evaluate MIST, a state-of-the-art deep learning approach, also based on fingerprint prediction. MIST assigns chemical subformulae to spectral peaks via energy-based modeling [22], then predicts a molecular fingerprint via a chemical formula-based transformer, and finally ranks the candidates by cosine similarity between the fingerprints [19]. Finally, we evaluate DeepSets [71]. The model processes spectra as sets of raw 2D peak representations, providing a complementary approach to FingerprintFFN and the state-of-the-art MIST which are based on alternative representations of spectra. DeepSets + Fourier features enhances DeepSets by using Fourier features enabling more accurate modeling of m/z values [20].
@@
[] Cosine Similarity $\uparrow$ Jensen-Shannon Similarity $\uparrow$ Hit Rate @ 1 $\uparrow$ Hit Rate @ 5 $\uparrow$ Hit Rate @ 20 $\uparrow$ Precursor m/z 0.15 (0.14-0.17) 0.59 (0.58-0.60) 0.38 (0.21-0.62) 1.72 (1.32-2.18) 7.17 (6.32-8.04) FFN Fingerprint 0.25 (0.24-0.26) 0.69 (0.63-0.65) 8.44 (7.56-9.34) 21.43 (20.10-22.79) 38.57 (36.99-40.23) GNN 0.19 (0.18-0.20) 0.64 (0.63-0.65) 3.95 (3.37-4.62) 11.92 (10.87-13.00) 26.27 (24.83-27.82) FraGNNet 0.52 (0.51-0.53) 0.91 (0.91-0.92) 46.64 (44.98-48.26) 72.56 (71.18-74.00) 83.58 (82.34-84.75) Bonus chemical formulae challenge Precursor m/z - - 2.09 (1.66-2.59) 8.52 (7.65-9.53) 22.65 (21.26-24.01) FFN Fingerprint - - 7.62 (6.77-8.54) 22.70 (21.32-24.12) 44.12 (42.51-45.75) GNN - - 3.63 (3.05-4.29) 13.55 (12.46-14.68) 33.77 (32.26-35.37) FraGNNet - - 31.93 (30.40-33.50) 63.20 (61.64-64.76) 82.70 (81.45-83.93)
@@ 5 Conclusions
This work was supported by the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (ID:90140) and by the Technology Agency of the Czech Republic through the project RETEMED (TN02000122). This work was also co-funded by the European Union (ERC FRONTIER No. 101097822; ELIAS No. 101120237). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. SH and AK are funded through Award R35GM148219 by the NIGMS of the National Institutes of Health. TP was supported by the Czech Science Foundation (GA CR) grant 21-11563M and by the European Union’s Horizon 2020 research and innovation programme under Marie Skłodowska-Curie grant agreement No. 891397. AY was supported by a Natural Sciences and Engineering Research Council of Canada (NSERC) scholarship and a Vector Institute research grant. LZ was supported by NSERC. FW was supported by Alberta Machine Intelligence Institute (AMII), and NSERC grant. RG was was supported by NSERC, AMII. DSW was supported by NSERC, and Genome Canada, Genome British Columbia and Genome Alberta. HR was supported by NSERC and the Canada Research Chair (CRC) Program. FK, KD and SB are supported by Deutsche Forschungsgemeinschaft (BO 1910/23). FK and SB are supported by the Ministry for Economics, Sciences and Digital Society of Thuringia (Framework ProDigital, DigLeben-5575/10-9). NAH and SB are supported by the Thüringer Ministerium für Wirtschaft, Wissenschaft und Digitale Gesellschaft (TMWWDG) with funds from the European Union as part of the European Regional Development Fund (ERDF, 2023 VFE 0003). JH and WB acknowledge support by the University of Antwerp Research Fund. FH was supported by Deutsche Forschungsgemeinschaft (DFG, 528775510). LL was supported by NSF Award 2239869. CB was supported by the Czech Academy of Sciences PPLZ fellowship number L200552251.